1. You decide to mate two flies and run analyses on the offspring. You mate a female fly that is true breeding for red eyes (R) and normal wings (N) with a male fly that is true breeding for black eyes (r) and vestigial wings (n). You then take one of the male fly babies from this cross, and mate it with a female fly that you have previously determined to have a genotype of RRNn. What Genotypic ratio would you expect? What Phenotypic ratio would you expect?

2. What is the probability of randomly picking a fly with the genotype RRNn from the offspring?

3. Jane is colorblind. She was worried that all her children would be colorblind as well. Luckily, her husband is not colorblind, so she assumed that her children would be all right. After all her ten children had grown to adulthood however, she was shocked to learn that ALL five of her sons were colorblind, while none of her daughters were. Even more puzzling, was that when her sons had children, none were colorblind, but all of her daughters had at least one son with the defect, but no daughters with it. Can you explain the inheritance pattern of colorblindness?
Answer #1

1. Eye color in Drosophila is X-linked.
Red eyes = XR = Dominant
White eyes = Xr = Recessive

Wing shape in Drosophila is an autosomal trait.
Normal wings = N = Dominant
Vestigial wings = n = Recessive

Parental cross: XRXR NN X XrY nn
F1: XRXrNn XRYNn

Next cross: XRYNn X XRXRNn
Gametes: (XRN) (XRn) (YN) (Yn) X (XRN) (XRn)
Progeny: (XRXRNN) (XRXRNn) (XRYNN) (XRYNn) (XRXRNn) (XRXRnn) (XRYNn) (XRYnn)
Genotypic ratio = 1:2:1:2:1:1
Phenotypic ratio:
Red normal females = 3
Red normal male = 3
Red vestigial female = 1
Red vestigial male = 1

2. The probability of RRNn = 2/8 = 1/4 = 0.25 = 25%

3. Colorblindenss is an X-linked trait.
Jane = XcXc
Jane's husband = XCY
Parental cross: XcXc X XCY
F1 progeny: XCXc XcY
All sons are affected and all daughters are carriers.
Daughters will give the Xc allele to their sons. So, male grandchildren will be affected.